On average, a painter misses 5 spots per room while painting houses. Write your answers to three decimal places.


What is the probability that in the next room he:


A) will forget to paint a maximum of 4 spots?


B) will forget no spots?


C) Find the mean of the number of unpainted spots in a house of 7 rooms.


1
Expert's answer
2021-06-15T12:39:20-0400

A) P(X≤4) = P(X=0) +P(X=1) +P(X=2) +P(X=3) +P(X=4)

"= e^{-λ}( \frac{λ^0}{0!} + \frac{λ^1}{1!}+\frac{λ^2}{2!}+\frac{λ^3}{3!}+\frac{λ^4}{4!}) \\

= e^{-5}( 1 + 5 + 12.5+20.83+26.04 ) \\

= 0.006737 \times 65.37 \\

=0.4404"

B) "P(X=0) = \frac{e^{-5} \times 5^0}{0!} = 0.006737"

C) The mean of the number of unpainted spots in a house of 7 rooms "= 7 \times 5 = 35"